# Renze Lou    Hanzi Xu    Sijia Wang    Jiangshu Du    Ryo Kamoi    Xiaoxin Lu    Jian Xie    Yuxuan Sun    Yusen Zhang    Jihyun Janice Ahn    Hongchao Fang    Zhuoyang Zou    Wenchao Ma    Xi Li    Kai Zhang    Congying Xia    Lifu Huang    Wenpeng Yin

### v2

@@ Abstract
Numerous studies have assessed the proficiency of AI systems, particularly large language models (LLMs), in facilitating everyday tasks such as email writing, question answering, and creative content generation. However, researchers face unique challenges and opportunities in leveraging LLMs for their own work, such as brainstorming research ideas, designing experiments, and writing or reviewing papers. In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in three fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not. AAAR-1.0 differs from prior benchmarks in two key ways: first, it is explicitly research-oriented, with tasks requiring deep domain expertise; second, it is researcher-oriented, mirroring the primary activities that researchers engage in on a daily basis. An evaluation of both open-source and closed-source LLMs reveals their potential as well as limitations in conducting sophisticated research tasks. We will keep iterating AAAR-1.0 to new versions. Project Webpage:   https://renzelou.github.io/AAAR-1.0/
@@ 1 Introduction
To this end, in this work, we introduce AAAR-1.0, a novel benchmark that aims to comprehensively assess the LLMs’ capacity on expert-level research tasks. As illustrated in Figure 1, AAAR-1.0 decomposes three distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews. To ensure data quality, senior AI researchers with extensive domain expertise perform data annotation for AAAR-1.0, followed by rigorous multi-round data examination and filtering. All three tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed. Crucially, tasks here are singular, stand-alone challenges (with clear input and output expectations) rather than a complicated task chain (Li et al., 2024; Lu et al., 2024), providing a more transparent assessment of the model’s intermediate output.
@@ 2 Related Work
With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024a; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024; Yu et al., 2024b), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024). For example, Si et al. (2024) conducted a large-scale human investigation on LLM-generated research ideas and found that LLMs can generate novel ideas compared with humans while lacking feasibility. Du et al. (2024) found that while LLMs are effective at summarizing papers, they tend to overly trust the authors’ claimed strengths and struggle to identify weaknesses specific to the paper. Furthermore, some works try to employ LLMs to solve more complicated research tasks that are composed of multiple steps (Li et al., 2024, 2023; Tang et al., 2023). Notably, Lu et al. (2024) proposed AI-Scientist, an autonomous agent framework that can handle a series of challenging research tasks consecutively, including generating research ideas, coming up with the corresponding experiments along with the implementations, and then writing the final research paper — exactly how human conduct a whole research pipeline. However, there is still a lack of systematic evaluations and quantitative analyses on the LLMs’ (intermediate) output of each single-step research task. Our work focuses on building a benchmark that has individual research steps with clear input-output expectations, thus making it suitable for comprehensive LLMs evaluation.
@@ 5.1 EquationInference
Table 1 shows the main results. Firstly, a simple baseline that predicts all equations as positive achieves 40% F1 (due to the 1:3 of positive and negative equations), while nearly all open-source LLMs even cannot beat this naive baseline. Notably, though the performance of Mixtral is slightly superior to the baseline, the extremely biased precision and recall scores imply that Mixtral is also simply predicting almost all samples as positive instead of truly inferring. Meanwhile, compared to the All-Positive baseline, the performance superiority of the strong close-source LLMs is not significant, the best LLM on this task only obtains 46.74%, which demonstrates the challenge of EqInfer compared with other similar benchmarks (Song et al., 2023). The generally high recall with low precision of all LLMs also indicates real-world risks, e.g., relying on LLMs to check the validity of equations in paper review.

### v4

@@ Abstract
Numerous studies have assessed the proficiency of AI systems, particularly large language models (LLMs), in facilitating everyday tasks such as email writing, question answering, and creative content generation. However, researchers face unique challenges and opportunities in leveraging LLMs for their own work, such as brainstorming research ideas, designing experiments, and writing or reviewing papers. In this study, we introduce AAAR-1.0, a benchmark dataset designed to evaluate LLM performance in four fundamental, expertise-intensive research tasks: (i) EquationInference, assessing the correctness of equations based on the contextual information in paper submissions; (ii) ExperimentDesign, designing experiments to validate research ideas and solutions; (iii) PaperWeakness, identifying weaknesses in paper submissions; and (iv) ReviewCritique, identifying each segment in human reviews is deficient or not. AAAR-1.0 differs from prior benchmarks in two key ways: first, it is explicitly research-oriented, with tasks requiring deep domain expertise; second, it is researcher-oriented, mirroring the primary activities that researchers engage in on a daily basis. An evaluation of both open-source and closed-source LLMs reveals their potential as well as limitations in conducting sophisticated research tasks. We will keep iterating AAAR-1.0 to new versions. Project Webpage:   https://renzelou.github.io/AAAR-1.0/
@@ 1 Introduction
To this end, in this work, we introduce AAAR-1.0, a novel benchmark that aims to comprehensively assess the LLMs’ capacity on expert-level research tasks. As illustrated in Figure 1, AAAR-1.0 decomposes four distinct expert-level AI research tasks from the researcher’s daily activities, including i) EquationInference, investigating whether the LLMs can infer the equation correctness based on the paper context; ii) ExperimentDesign, validating LLMs’ ability on designing reliable experiments for a research idea; iii) PaperWeakness, testing the quality of weaknesses discovered by LLMs from paper drafts; and iv) ReviewCritique, investigating whether LLMs can identify and explain the deficient/unreliable human-written paper reviews. To ensure data quality, senior AI researchers with extensive domain expertise perform data annotation for AAAR-1.0, followed by rigorous multi-round data examination and filtering. All tasks require models to possess strong domain knowledge covering various cutting-edge research findings, as well as expert-level research experience, to the extent that even humans need substantial research accumulation to tackle the tasks we designed. Crucially, tasks here are singular, stand-alone challenges (with clear input and output expectations) rather than a complicated task chain (Li et al., 2024; Lu et al., 2024), providing a more transparent assessment of the model’s intermediate output.
@@ 2 Related Work
With the rapid evolution of pertaining techniques, LLMs are found to be useful in assisting various research disciplines (Yu et al., 2024a; Labrak et al., 2024), particularly in AI research, such as generating novel research ideas (Kumar et al., 2024; Yu et al., 2024b), reviewing research draft (Gao et al., 2024; Du et al., 2024; Liang et al., 2024; Zhu et al., 2025), and writing scientific papers (Chamoun et al., 2024; Lu et al., 2024; Weng et al., 2024). For example, Si et al. (2024) conducted a large-scale human investigation on LLM-generated research ideas and found that LLMs can generate novel ideas compared with humans while lacking feasibility. Du et al. (2024) found that while LLMs are effective at summarizing papers, they tend to overly trust the authors’ claimed strengths and struggle to identify weaknesses specific to the paper. Furthermore, some works try to employ LLMs to solve more complicated research tasks that are composed of multiple steps (Li et al., 2024, 2023; Tang et al., 2023). Notably, Lu et al. (2024) proposed AI-Scientist, an autonomous agent framework that can handle a series of challenging research tasks consecutively, including generating research ideas, coming up with the corresponding experiments along with the implementations, and then writing the final research paper — exactly how human conduct a whole research pipeline. However, there is still a lack of systematic evaluations and quantitative analyses on the LLMs’ (intermediate) output of each single-step research task. Accordingly, our work focuses on building a benchmark consisting of individual research steps with clear input-output expectations, making it suitable for comprehensive LLM evaluation. Moreover, we emphasize that relying on LLMs to fully replace human effort might compromise academic integrity. While our benchmark primarily serves an educational purpose — LLMs assist junior researchers by providing imperfect but insightful ideas, rather than by governing the entire research process.
@@ 5.1 EquationInference
Table 1 shows the main results. Firstly, a simple baseline that predicts all equations as positive achieves 40% F1 (due to the 1:3 of positive and negative equations), while nearly all open-source LLMs even cannot beat this naive baseline. Notably, though the performance of Mixtral is slightly superior to the baseline, the extremely biased precision and recall scores imply that Mixtral is also simply predicting almost all samples as positive instead of truly inferring. Meanwhile, compared to the All-Positive baseline, the performance superiority of the strong close-source LLMs is not significant, the best LLM on this task only obtains 47.98%, which demonstrates the challenge of EqInfer compared with other similar benchmarks (Song et al., 2023). The generally high recall with low precision of all LLMs also indicates real-world risks, e.g., relying on LLMs to check the validity of equations in paper review.
@@ 6 Conclusion
Limitations
We shed light on two limitations of this work: i) As we gather data from open-source platforms such as arXiv and OpenReview, there is a possibility that current or future LLMs may be trained on the same source data utilized in our benchmark. This situation could influence the fairness of LLM comparisons and the conclusions drawn from this paper. While we acknowledge this potential data leakage, we maintain that our work can provide valuable insights and serve as an upper bound for some LLMs, particularly if they are indeed pretrained on those scientific papers. At present, this research can inform the design of current LLMs and agents, enhancing the community’s understanding of the strengths and limitations of using LLMs in scientific research. ii) Meanwhile, due to the high consumption of employing senior researchers in conducting data annotation, the data size for some tasks, such as ExpDesign, is relatively small. This might introduce variability in LLM performance, even with repeated runs. As this work marks the beginning of a series, we plan to release larger-scale datasets that will cover more diverse research tasks in AAAR-2.0.
Acknowledgments
We would like to thank arXiv and OpenReview for releasing the paper source packages and the review comments. The authors would also like to thank Ibraheem Moosa and Sarkar Snigdha Sarathi Das for assisting in the data collection.